GlueX TOF Meeting, November 3, 2015

From GlueXWiki
Jump to: navigation, search

GlueX Time-of-Flight Meeting
Tuesday, November 3, 2015
11:00 am EST
JLab: CEBAF Center, Room F326/327


  1. Announcements
  2. Review of minutes from September 22
  3. Calibration Status FSU
  4. Calibration Status JLab
  5. NIM Article



Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at .



  • FSU: Hussein Al Ghoul, Brad Cannon, Sasha Ostrovidov, Aristeidis Tsaris
  • JLab: Mark Ito (chair), Simon Taylor

There is a recording of this meeting at .

Calibration with Cosmics

Sasha gave a report on recent work.

  • He has succeeded in taking cosmic data, but had a couple of difficulties:
    1. He needed to attempt to start runs several times before getting any triggers at all. He suspects that with multi-block readout of the CAEN TDCs, the token passing between modules would stop, hanging the system.
    2. ROCTOF1, which houses the FADCs, would crash in minutes from the start of the run. A hard reboot was needed to get things going again.
      • This problem has resolved itself recently.
    • In the end he got a few times 100 k events.
  • The trend of higher gains at the edges of the arrays seen in beam data is also seen in cosmic data.
    • This rules out the magnetic field, or a systematic problem using beam tracks.
    • Possible culprits:
      • Different HV supplies than those used at FSU.
      • Different HV cables.
    • He is working on a program to adjust HV to equalize gains. Data will be taken, gains measured, HV adjusted, and process iterated.
    • He is also seeing a difference in the gain when comparing a hit in the middle of the paddle versus one near the ends, even after accounting for attenuation. The gain is about 10% higher near the ends.
  • He is trying to judge the effect of gain calibration using a set of "tracks" in the TOF in cosmic data. He requires 5 counters minimum. The yield of such events is extremely low, about one part in ten thousand.
    • The trigger rate is about 600 Hz.
    • The threshold for a single FADC channel to be read out is 500 counts. The trigger threshold on the sum is about 50,000 typically, but other values have been tried.


  • Brad is in contact with Sean Dobbs about a plug-in to be used in the calibration challenge.
  • He is also cleaning up the script to drive the calibrations already done.